A kid has 3 times more money than another . If the first one spends 2/7 of his sum of money and the second one spends 3/5 of his sum of money.So, the first kid will have 122 $ more than the second one.Find the sum of money that each had at the begining .
Answer by checkley71(8403)

LET X=THE AMOOUNT OF MONEY FOR THE FIRST KID.
LET 3X=THE AMOUNT OF MONEY FOR THE SECOND KID.
IF THE FIRST KID SPENDS 2/7 THEN HE HAS 5*3X/7 LEFT.
THE SECOND KIS SPENDS 3/5 THEN HE HAS 2X/5 LEFT.
5*3x/7=2x/5+122
15x/7-2x/5=122
(5*15x)+7*-2x)/7*5=122
75x-14x/35=122
61x=35*122
61x=4270
x=4270/61
x=70 for the amount of the second kid.
3*70=210 for the first kid.
PROOF:
15*70/7=2*70/5+122
15*10=140/5+122
150=28+122
150=150